Calvin paid his $1.35 skate rental with dimes and nickels only. There were 19 coins in all. How many of each coin were there?
Answer by checkley71(8403)

N+D=19 N=19-D
.10D+.05(19-D)=1.35
.10D+.95-.05D=1.35
.05D=1.35-.95
.05D=.40
D=.40/.05
D=8 DIMES.
19-8=11 NICKELS.
PROOF
.10*8+.05*11=1.35
.80+.55=1.35
1.35=1.35